OpenStudy (anonymous):

WILL MEDAL AND FAN BEST ANSWER!! I NEED HELP!! I need the answer AND an explanation Subtract. 1 over 2 - 2 over 5

OpenStudy (anonymous):

1/10 This is very simple. If you have a calculator or can add,subtract, multiply and divide fractions. But nonetheless that is your answer:D

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

i don't have a fraction calculator.

OpenStudy (anonymous):

oh well their is your answer

OpenStudy (kropot72):

First you need to find a common denominator for both the fractions. This can be done by multiplying the denominators: 2 * 5 = 10. So next we will make both of the fractions tenths, as follows: \[\large \frac{1}{2}\times\frac{5}{5}=\frac{5}{10}\] \[\large \frac{2}{5}\times\frac{2}{2}=\frac{4}{10}\] Now the calculation becomes \[\large \frac{5}{10}-\frac{4}{10}=\frac{1}{10}\]

OpenStudy (anonymous):

@kropot72